OpenStudy (anonymous):

A cell phone company found that 15 of their phones sold in June were defective. The company sold 250 phones in June. How many of the 300 phones should they expect to be defective?

OpenStudy (tcp2315):

do you mean the next month they made 300, and how many might be defective?

OpenStudy (tcp2315):

if yes then its simply x/300 = 15/250 x= (300 x 15)/250 x= 18

OpenStudy (anonymous):

15 divided by 250 = 0.06 or 6%... 6% of 300 = ? or 0.06 of ( times ) 300 = ?

OpenStudy (anonymous):

and TCP2315 gave the whole game away :-)